• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 02:58:26 on November22,2009

    Entry number 301720

    keyword=Compton running conditions

    The Compton is currently running at HV=-1400 and completely out of the beam. We will run like this over the night, unless something becomes unstable, in which case we will lower the HV.

    The laser is also not locking properly.

    At HV=-1500, there are photon detector base current fluctuations with beam trips (hence we're running at 1400). Rates were very high a few minutes ago, and have since become normal. I assume this is due to vacuum problems, but it also could be due to the shielding changes that were made earlier today.


    FIGURE 1

    FIGURE 2